OpenStudy (anonymous):

Asegment has endpoints at (-13, 19) and (4, -7). What is the midpoint of the segment?

OpenStudy (anonymous):

Midpoint formula is \[\frac{ x1+x2 }{ 2 }; \frac{ y1+y2 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ -13+4 }{ 2 }; \frac{ 19+(-7) }{ 2 }\]